# Distribution-free data-driven smooth tests without $\chi^{2}$

### v2

@@ 3 Smooth tests via projected bootstrap
The classical parametric bootstrap requires re-estimating the unknown parameter $\bm{\beta}$ on each bootstrap sample to account for the variability introduced by the estimation of the parameters. In some instances, however, a repeated estimation can make the procedure computationally intensive. The so-called “projected bootstrap” (Algeri,, 2022; Algeri and Khmaladze,, 2024) overcomes this limitation by exploiting the projection structure induced by parameter estimation – a result first established by Khmaladze, (1980) – to avoid repeating the estimation of the parameter on each bootstrap replicate. As shown in what follows, such a projection arises in the context of smooth tests rather organically.
Let the true (but unknown) parameter vector be $\bm{\beta}_{0}$, that is, under the null hypothesis $H_{0},$ the true distribution $Q$ equals $G_{\bm{\beta}_{0}}$. Let $\mathcal{N}_{\bm{\beta}_{0}}\subset\mathcal{B}$ denote the closure of a given neighborhood of $\bm{\beta}_{0}$ and denote with $\mathcal{L}\left(G_{\bm{\beta}}\right)$ the subspace of $L^{2}\left(G_{\bm{\beta}}\right)$ given by
$\mathcal{L}\left(G_{\bm{\beta}}\right)=\left\{h_{\bm{\beta}}\in L^{2}\left(G_{\bm{\beta}}\right):\left\langle h_{\bm{\beta}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=0\right\}.$
We make the following regularity assumptions:
@@
    The estimator $\widehat{\bm{\beta}}_{n}$ has the following asymptotic representation
    $\sqrt{n}(\widehat{\bm{\beta}}_{n}-\bm{\beta}_{0})=v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}})+o_{P}(1),$
    where $\bm{\psi}_{\bm{\beta}}$ is a $p$-dimensional vector function in $\mathcal{L}\left(G_{\bm{\beta}}\right)$ and is continuously differentiable with respect to $\bm{\beta}$.
  2. (A2)
    For any $h_{\bm{\beta}}\in\mathcal{L}\left(G_{\bm{\beta}}\right)$, its gradient taken with respect to $\bm{\beta}$, denoted by $\nabla_{\bm{\beta}}h_{\bm{\beta}}$, exists and is uniformly continuous for all $x\in\mathcal{X}$ and $\bm{\beta}\in\mathcal{N}_{\bm{\beta}_{0}}$.
@@
    $\int_{\mathcal{X}}\nabla_{\bm{\beta}}\Big{[}h_{\bm{\beta}}(x)dG_{\bm{\beta}}(x)\Big{]}=\nabla_{\bm{\beta}}\int_{\mathcal{X}}h_{\bm{\beta}}(x)dG_{\bm{\beta}}(x).$
To identify the projection structure arising from parameter estimation, we begin by demonstrating that the process $v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})$ is asymptotically equivalent to its first-order Taylor expansion around the true parameter $\bm{\beta}_{0}$.
@@
Let $r_{n}$ be defined by the equation
$v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})=v_{G,n}(h_{\bm{\beta}_{0}})+v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}+r_{n}.$ (10)
If the assumptions (A1)-(A2) are satisfied, then $r_{n}=o_{P}(1)$.
@@
The result can be demonstrated by expressing $r_{n}$ as follows:
$r_{n}$ $=\sqrt{n}(\bm{\widehat{\beta}}_{n}-\bm{\beta}_{0})^{T}\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}^{{}^{\prime}}}-v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}$
$=\sqrt{n}(\bm{\widehat{\beta}}_{n}-\bm{\beta}_{0})^{T}\frac{1}{\sqrt{n}}\left[\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}^{{}^{\prime}}}-\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}\right]$
$\quad\quad+\Big{[}\sqrt{n}(\bm{\widehat{\beta}}_{n}-\bm{\beta}_{0})^{T}-v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\Big{]}\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}},$
where $\bm{\beta}$$^{{}^{\prime}}$ lies between $\bm{\beta}_{0}$ and $\widehat{\bm{\beta}}_{n}$. Under the assumptions (A1)-(A2) and by the law of large numbers, it follows that $r_{n}=o_{P}(1)$ as $n\to\infty.$
The leading terms of the right-hand side of the asymptotic representation in (10) can also be expressed in function-parametric form; as shown in Proposition 2, the functions indexing the process in (2) are projections of the functions $h_{\bm{\beta}_{0}}$.
@@
If the assumptions (A1)-(A3) hold, then
$v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})=v_{G,n}(\Pi h_{\bm{\beta}_{0}})+o_{P}(1),$ (11)
where
$\Pi h_{\bm{\beta}}=h_{\bm{\beta}}-\bm{\psi}_{\bm{\beta}}^{T}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}$ (12)
is a projection of $h_{\bm{\beta}}$ and it is orthogonal to $\bm{u}_{\bm{\beta}}$ when $\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}=I_{p}$.
Proof.
Assumption (A3) and the law of large numbers imply
$\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})$ $=\frac{1}{\sqrt{n}}\int_{\mathcal{X}}\nabla_{\bm{\beta}}\left[h_{\bm{\beta}}(x)dv_{\bm{\beta},n}(x)\right]$ (13)
$=\frac{1}{\sqrt{n}}\int_{\mathcal{X}}[\nabla_{\bm{\beta}}h_{\bm{\beta}}(x)]dv_{\bm{\beta},n}(x)-\int_{\mathcal{X}}h_{\bm{\beta}}(x)\nabla_{\bm{\beta}}g_{\bm{\beta}}(x)dx\overset{p}{\longrightarrow}-\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}.$
Substituting (13) into (10) gives the results in (11). Additionally, when $\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}=I_{p}$,
$\Pi^{2}h_{\bm{\beta}}=h_{\bm{\beta}}-2\bm{\psi}_{\bm{\beta}}^{T}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}+\bm{\psi}_{\bm{\beta}}^{T}\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\Pi h_{\bm{\beta}},$
that is, $\Pi$ is a projection, and since
$\left\langle\bm{u}^{T}_{\bm{\beta}},\Pi h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}-\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=0,$
such a projection is orthogonal to $\bm{u}_{\bm{\beta}}.$
The result in equations (11)-(12) can be equivalently stated as
$v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})=v_{G,n}(h_{\bm{\beta}_{0}})-v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\left\langle\bm{u}^{T}_{\bm{\beta}_{0}},h_{\bm{\beta}_{0}}\right\rangle_{G_{\bm{\beta}_{0}}}+o_{P}(1),$
by virtue of the linearity of the projector $\Pi$. Moreover, the condition $\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}=I_{p}$ is satisfied for many standard estimators, such as the maximum likelihood and the method of moments estimators. For instance, when the parameters are estimated via maximum likelihood, we have
$\bm{\psi}_{\bm{\beta}}(x)=\Gamma^{-1}_{\bm{\beta}}\bm{u}_{\bm{\beta}}(x),\quad\text{ and }\quad\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{u}_{\bm{\beta}}^{T}\right\rangle_{G_{\bm{\beta}}}\Gamma^{-1}_{\bm{\beta}}=I_{p}.$
Furthermore, let $\mu_{k,\bm{\beta}}$ be the $k$-thth moment of $X$ under $G_{\bm{\beta}}$ and let $\bm{\varphi}_{\bm{\beta}}(x)$ be a $p$-dimensional vector-valued function with elements $x^{k}-\mu_{k,\bm{\beta}}$, for $k=1,\dots,p$. If $\widehat{\bm{\beta}}_{n}$ is the method of moments estimator, it solves the estimating equations:
$v_{G,n}(\bm{\varphi}_{\widehat{\bm{\beta}}_{n}})=0,$ (14)
and the first-order Taylor expansion of (14) around $\widehat{\bm{\beta}}_{n}=\bm{\beta}_{0}$ leads to
$\sqrt{n}(\widehat{\bm{\beta}}_{n}-\bm{\beta}_{0})=-\frac{1}{\sqrt{n}}\left[\nabla_{\bm{\beta}}v_{G,n}(\bm{\varphi}_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}\right]^{-1}v_{G,n}(\bm{\varphi}_{\bm{\beta}_{0}})+o_{P}(1),$
from which, following a similar argument as in (13), we obtain $\nabla_{\bm{\beta}}v_{G,n}(\bm{\varphi}_{\bm{\beta}})\overset{p}{\longrightarrow}-\big{\langle}{\bm{\varphi}^{T}_{\bm{\beta}},\bm{u}^{T}_{\bm{\beta}}\big{\rangle}}_{G_{\bm{\beta}}}$. Therefore,
$\bm{\psi}_{\bm{\beta}}(x)=\bigl{\langle}{\bm{\varphi}^{T}_{\bm{\beta}},\bm{u}^{T}_{\bm{\beta}}\bigl{\rangle}}_{G_{\bm{\beta}}}^{-1}\bm{\varphi}_{\bm{\beta}}(x)\quad\text{ and }\quad\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{\psi}^{T}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{\varphi}^{T}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}\left[\left\langle\bm{\varphi}^{T}_{\bm{\beta}},\bm{u}^{T}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}^{T}\right]^{-1}=I_{p}.$
In the context of smooth tests, Proposition 2 implies that the process $v_{G,n}$ indexed by the functions $\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}$ and that indexed by functions $\Pi\widetilde{h}_{j\bm{\beta}_{0}}$ are asymptotically equal. Furthermore, the functions $\widetilde{h}_{j\bm{\beta}}$ are orthogonal to $\bm{u}_{\bm{\beta}}$ by the definition in equation (3). It follows that
$\Pi\widetilde{h}_{j\bm{\beta}_{0}}=\widetilde{h}_{j\bm{\beta}_{0}}-\bm{\psi}_{\bm{\beta}_{0}}^{T}\bigl{\langle}{\bm{u}^{T}_{\bm{\beta}_{0}},\widetilde{h}_{j\bm{\beta}_{0}}\bigl{\rangle}}_{G_{\bm{\beta}_{0}}}=\widetilde{h}_{j\bm{\beta}_{0}};$
therefore, the asymptotic equality in (11) can be further simplified as in Proposition 3.
Proposition 3.
Under condition (A1), and assuming that the functions $\widetilde{h}_{j\bm{\beta}}$ satisfy the assumptions (A2) and (A3) for each $j\in\mathcal{M}_{n}$, we have
$v_{G,n}(\widetilde{h}_{j\bm{\beta}_{0}})=v_{G,n}(\Pi\widetilde{h}_{j\bm{\beta}_{0}})=v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})+o_{P}(1).$ (15)
@@
To gain some insights on the computational advantages entailed by Proposition 3, denote with $\widehat{\bm{\beta}}_{\text{obs}}$ the estimate of $\bm{\beta}$ obtained on the observed data. In the parametric bootstrap, such an estimate plays the same role as $\bm{\beta}_{0}$ in (15). Let $\widehat{\bm{\beta}}_{n}^{(b)}$ be the parameter estimate obtained on the $b$-th bootstrap sample generated from $G_{\widehat{\bm{\beta}}_{\text{obs}}}.$ Equation (15) implies that the empirical process indexed by functions $\widetilde{h}_{j\widehat{\bm{\beta}}^{(b)}_{n}}$ is asymptotically equal to the process indexed by $\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}}$. In other words, the projection structure characterizing the functions $\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}}$ makes the effect of re-estimating the parameter asymptotically negligible. Thus, for sufficiently large $n$, one only needs to evaluate the process $v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}})$ over different bootstrap samples but needs not to repeat the estimation of $\widehat{\bm{\beta}}_{n}^{(b)}$ at each replicate. It follows that this projected bootstrap procedure can always make the simulation more efficient. It is particularly advantageous when dealing with computationally intensive postulated models and when parameter estimation is time-consuming. Moreover, the consistency of the projected bootstrap in recovering the true distribution of the function-parametric empirical process $v_{G,n}(\widetilde{h}_{j\bm{\beta}_{0}})$ is guaranteed under the same conditions needed for the classical parametric bootstrap (cfr. Babu and Rao,, 2004).
3.1 Empirical illustrations of smooth tests via projection

### v3

@@ 3 Smooth tests via projected bootstrap
The classical parametric bootstrap requires re-estimating the unknown parameter $\bm{\beta}$ on each bootstrap sample to account for the variability introduced by estimation. In some instances, however, a repeated estimation can make the procedure computationally intensive. The so-called “projected bootstrap” (Algeri, 2022; Algeri and Khmaladze, 2026) overcomes this limitation by exploiting the projection structure induced by parameter estimation – a result first established by Khmaladze (1980) – to avoid repeating the estimation of the parameter on each bootstrap replicate. As shown in what follows, such a projection arises in the context of smooth tests rather organically.
Let the true (but unknown) parameter vector under the null be $\bm{\beta}_{0}$, that is, if $H_{0}$ in (2) holds, then $Q\equiv G_{\bm{\beta}_{0}}$. Denote with $\mathcal{L}\left(G_{\bm{\beta}}\right)$ the subspace of $L^{2}\left(G_{\bm{\beta}}\right)$ given by
$\mathcal{L}\left(G_{\bm{\beta}}\right)=\left\{h\in L^{2}\left(G_{\bm{\beta}}\right):\left\langle h,\bm{1}\right\rangle_{G_{\bm{\beta}}}=0\right\}.$
In what follows, we focus our attention on the subset of functions in $\mathcal{L}\left(G_{\bm{\beta}}\right)$ that depend on $\bm{\beta}$. Such functions primarily intervene in our study by either characterizing the alternative in (1) or defining the estimating equations for $\bm{\beta}$. We assume such functions to be continuously differentiable in $\bm{\beta}$ and satisfy the usual regularity conditions used in standard asymptotic arguments (cf. Van der Vaart, 2000, Ch.5) and which imply the following:
@@
    Let $\widehat{\bm{\beta}}_{n}$ be a consistent root of the system of estimating equations
    $v_{\widehat{\bm{\beta}}_{n},n}(\bm{\psi}_{\widehat{\bm{\beta}}_{n}})=0,$ (10)
    where $\bm{\psi}_{\bm{\beta}}$ is a $p$-dimensional vector function with linearly independent components in $\mathcal{L}\left(G_{\bm{\beta}}\right)$. Denote with $\nabla_{\bm{\beta}}h_{\bm{\beta}}$ the gradient of $h_{\bm{\beta}}$ taken with respect to $\bm{\beta}$. The Taylor expansion
    $v_{\widehat{\bm{\beta}}_{n},n}(h_{\widehat{\bm{\beta}}_{n}})=v_{\bm{\beta}_{0},n}(h_{\bm{\beta}_{0}})+(\widehat{\bm{\beta}}_{n}-\bm{\beta}_{0})^{T}\nabla_{\bm{\beta}}v_{\bm{\beta},n}(h_{\bm{\beta}})\Bigl|_{\bm{\beta}=\bm{\beta}_{0}}+o_{P}(1)$ (11)
    is valid.
@@
    $\nabla_{\bm{\beta}}\int_{\mathcal{X}}h_{\bm{\beta}}(x)dG_{\bm{\beta}}(x)=\int_{\mathcal{X}}\nabla_{\bm{\beta}}\Big[h_{\bm{\beta}}(x)dG_{\bm{\beta}}(x)\Big].$
To ease the intuition, in what follows, we assume (A1)-(A2) directly.
As shown in the proposition below, when replacing $\bm{\beta}$ with $\widehat{\bm{\beta}}_{n}$ in $v_{{\bm{\beta}},n}(h_{{\bm{\beta}}})$ the resulting process is asymptotically equal to a projection of the latter evaluated at the true parameter value $\bm{\beta}_{0}$.
@@
If (A1)-(A2) hold, then, under $H_{0},$
$v_{\widehat{\bm{\beta}}_{n},n}(h_{\widehat{\bm{\beta}}_{n}})=v_{\bm{\beta}_{0},n}(\Pi h_{\bm{\beta}_{0}})+o_{P}(1),$ (12)
where
$\Pi h_{\bm{\beta}}=h_{\bm{\beta}}-\bm{\psi}_{\bm{\beta}}^{T}\left\langle\bm{u}_{\bm{\beta}},\bm{\psi}^{T}_{\bm{\beta}}\right\rangle^{-1}_{G_{\bm{\beta}}}\left\langle\bm{u}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}$ (13)
is a projection of $h_{\bm{\beta}}$.
@@
Under Assumption (A2), we have,
$\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{\bm{\beta},n}(h_{\bm{\beta}})\Big|_{\bm{\beta}=\bm{\beta}_{0}}$ $=\frac{1}{\sqrt{n}}\int_{\mathcal{X}}\nabla_{\bm{\beta}}\!\left[h_{\bm{\beta}}(x)\,dv_{\bm{\beta},n}(x)\right]\Big|_{\bm{\beta}=\bm{\beta}_{0}}$ (14)
$=\frac{1}{\sqrt{n}}\int_{\mathcal{X}}\left[\nabla_{\bm{\beta}}h_{\bm{\beta}}(x)\right]\Big|_{\bm{\beta}=\bm{\beta}_{0}}\,dv_{\bm{\beta}_{0},n}(x)-\int_{\mathcal{X}}h_{\bm{\beta}_{0}}(x)\,\left[\nabla_{\bm{\beta}}g_{\bm{\beta}}(x)\right]\Big|_{\bm{\beta}=\bm{\beta}_{0}}\,dx$
$=-\,\left\langle\bm{u}_{\bm{\beta}_{0}},h_{\bm{\beta}_{0}}\right\rangle_{G_{\bm{\beta}_{0}}}+o_{P}(1),$
in which the last equality holds by the law of large numbers. Substituting the right-hand side of such an equality into (11) gives:
$v_{\widehat{\bm{\beta}}_{n},n}(h_{\widehat{\bm{\beta}}_{n}})=v_{\bm{\beta}_{0},n}(h_{\bm{\beta}_{0}})-\sqrt{n}(\widehat{\bm{\beta}}_{n}-\bm{\beta}_{0})^{T}\left\langle\bm{u}_{\bm{\beta}_{0}},h_{\bm{\beta}_{0}}\right\rangle_{G_{\bm{\beta}_{0}}}+o_{P}(1).$
Since the components of $\bm{\psi}_{{\bm{\beta}}}$ belong to $\mathcal{L}\left(G_{\bm{\beta}}\right)$, the asymptotic equality above also holds for $v_{\widehat{\bm{\beta}}_{n},n}(\bm{\psi}_{\widehat{\bm{\beta}}_{n}})$ and equating it to zero leads to:
$\sqrt{n}(\widehat{\bm{\beta}}_{n}-\bm{\beta}_{0})=\left\langle\bm{u}_{\bm{\beta}_{0}},\bm{\psi}^{T}_{\bm{\beta}_{0}}\right\rangle^{-1}_{G_{\bm{\beta}_{0}}}v_{\bm{\beta}_{0},n}(\bm{\psi}_{\bm{\beta}_{0}})+o_{P}(1).$
Combining the last two expressions yields:
$v_{\widehat{\bm{\beta}}_{n},n}(h_{\widehat{\bm{\beta}}_{n}})=v_{\bm{\beta}_{0},n}(h_{\bm{\beta}_{0}})-v_{\bm{\beta}_{0},n}(\bm{\psi}^{T}_{\bm{\beta}_{0}})\left\langle\bm{u}_{\bm{\beta}_{0}},\bm{\psi}^{T}_{\bm{\beta}_{0}}\right\rangle^{-1}_{G_{\bm{\beta}_{0}}}\left\langle\bm{u}_{\bm{\beta}_{0}},h_{\bm{\beta}_{0}}\right\rangle_{G_{\bm{\beta}_{0}}}+o_{P}(1),$
which can be equivalently expressed as in (12)-(13) by linearity of $\Pi$. Additionally, it is easy to verify that $\Pi\Pi h_{\bm{\beta}}=\Pi h_{\bm{\beta}},$ thus, $\Pi$ is a projection operator. ∎
In general, the projection given by $\Pi$ is not orthogonal; that is because
$\Pi\bm{\psi}_{\bm{\beta}}=0\quad\text{and}\quad\left\langle\bm{u}_{\bm{\beta}},\Pi h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=0$
while neither $\Pi\bm{u}_{\bm{\beta}}$ nor $\left\langle\bm{\psi}_{\bm{\beta}},\Pi h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}$ are zero. From the above expressions, however, it is clear that $\Pi$ defines an orthogonal projection when $\bm{\psi}_{\bm{\beta}}\equiv\bm{u}_{\bm{\beta}}$, that is, when the estimation is conducted via maximum likelihood.
In the context of smooth tests, Proposition 1 implies that the processes $v_{\widehat{\bm{\beta}}_{n},n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})$ and $v_{{\bm{\beta}}_{0},n}(\Pi\widetilde{h}_{j\bm{\beta}_{0}})$ are asymptotically equal. Furthermore, the functions $\widetilde{h}_{j\bm{\beta}}$ are orthogonal to $\bm{u}_{\bm{\beta}}$ by construction (see Equation (3)). It follows that
$\Pi\widetilde{h}_{j\bm{\beta}_{0}}=\widetilde{h}_{j\bm{\beta}_{0}}-\bm{\psi}_{\bm{\beta}_{0}}^{T}\left\langle\bm{u}_{\bm{\beta}_{0}},\bm{\psi}^{T}_{\bm{\beta}_{0}}\right\rangle^{-1}_{G_{\bm{\beta}_{0}}}\bigl\langle{\bm{u}_{\bm{\beta}_{0}},\widetilde{h}_{j\bm{\beta}_{0}}\bigl\rangle}_{G_{\bm{\beta}_{0}}}=\widetilde{h}_{j\bm{\beta}_{0}}.$
This fact, combined with the asymptotic equality in (12), brings us to the following proposition.
@@
Assume that (A1) and (A2) hold for each $\widetilde{h}_{j\bm{\beta}}$, $j\in\mathcal{M}_{n}$, then
$v_{\bm{\beta}_{0},n}(\widetilde{h}_{j\bm{\beta}_{0}})=v_{\bm{\beta}_{0},n}(\Pi\widetilde{h}_{j\bm{\beta}_{0}})=v_{\widehat{\bm{\beta}}_{n},n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})+o_{P}(1).$ (15)
@@
To understand the computational advantages entailed by Proposition 2, denote with $\widehat{\bm{\beta}}_{\text{obs}}$ the estimate of $\bm{\beta}$ obtained on the observed data by solving (10). In the parametric bootstrap, such an estimate plays the same role as $\bm{\beta}_{0}$ in (15). Let $\widehat{\bm{\beta}}_{n}^{(b)}$ be the parameter estimate obtained on the $b$-th bootstrap sample generated from $G_{\widehat{\bm{\beta}}_{\text{obs}}}.$ Equation (15) implies that the empirical process indexed by functions $\widetilde{h}_{j\widehat{\bm{\beta}}^{(b)}_{n}}$ is asymptotically equal to the process indexed by $\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}}$. In other words, the projection structure characterizing the functions $\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}}$ makes the effect of re-estimating the parameter asymptotically negligible. Thus, for sufficiently large $n$, the user must evaluate the process $v_{\widehat{\bm{\beta}}_{\text{obs}},n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}})$ over different bootstrap samples but does not need to repeat the estimation of $\bm{\beta}$ at each replicate. It follows that this projected bootstrap procedure can always make the simulation more efficient. As demonstrated in the next section with an example, the projected bootstrap is particularly advantageous when parameter estimation is time-consuming.
Finally, the consistency of the projected bootstrap in recovering the true distribution of the function-parametric empirical process $v_{\bm{\beta}_{0},n}(\widetilde{h}_{j\bm{\beta}_{0}})$ is guaranteed under the same conditions needed for the classical parametric bootstrap (cfr. Babu and Rao, 2004).
3.1 Empirical illustrations of smooth tests via the projected bootstrap